Find the solution of the initial-value problem and determine its interval of existence (i.e domain of the resulting function y).

(i)dy/dt=(1+t²)/t , y(t = 1) = 0

(ii) (t + 1) dy/dt = 1- y , y(t = 0) = 3 [Hint: Let A = -(±e^(-c))]
1
Expert's answer
2020-06-29T19:16:39-0400

(i) "\frac{dy}{dt}=\frac{1+t^2}{t} , y(t = 1) = 0."

We can separate variables and integrate this equation:

"dy=\frac{1+t^2}{t}dt,"

"dy=(\frac{1}{t}+t)dt,"

"\int dy = \int (\frac{1}{t}+t)dt,"

"y(t)=ln|t|+\frac{t^2}{2}+C," "C" is a constant.

Next, we plug in our initial condition: "y(t=1)=0=ln|1|+\frac{1^2}{2}+C=0.5+C."

We get that "C=-0.5."

Thus, plugging back into our solution for y, we get:

"y(t)=ln|t|+\frac{t^2}{2}-0.5."

It's domain is "t\neq 0" or "t \in (-\infty;0)\cup(0;+\infty)" .


(ii) "(t+1)\frac{dy}{dt} = 1- y , y(t = 0) = 3."

We can separate and integrate:

"\frac{dy}{1-y}=\frac{dt}{t+1},"

"-\frac{d(1-y)}{1-y}=\frac{d(1+t)}{dt},"

"\u2212ln\u22231\u2212y\u2223=ln\u22231+t\u2223+C," "C" is a constant.

"ln|(1-y)(1+t)|=-C,"

"(1-y)(1+t)=e^{-C}."

Set "A=e^{-C}," hence we have:

"(1\u2212y)(1+t)=A," where A is a constant.

"1-y=\frac{A}{1+t},"

"y=1-\frac{A}{1+t}=\frac{1+t-A}{1+t}=\frac{t+C_1}{t+1}," where "C_1=1-A" is a constant.

Next, we plug in our initial condition:

"y(t=0)=3=\frac{0+C_1}{0+1}=C_1," and we get that "C_1=3."

Thus, plugging back into our solution for y, we get:

"y(t)=\frac{t+3}{t+1}," it's domain is "t \neq -1" or "t \in (-\infty; -1)\cup (-1; +\infty)."